# Progressive Checkerboards for Autoregressive Multiscale Image Generation

### v1

@@ 4.3 Relationship Between Scale Ratio and Sampling Steps
We explore the performance of our model for different multiscale scale factors and inference steps. Using the small (S) sized model, we train models using four scale-up factors: $\sqrt{2},2,3$ and $4$, as well as a single-scale baseline. The exact sizes are shown in Fig 4(b). We then evaluate each model using different numbers of sampling steps, varying the number of steps per scale, which corresponds to the partitioning size of the checkerboard pattern. Note that in our implementation, each scale ratio requires retraining, but the number of sampling steps can be varied at inference.
We evaluate each point using FID and IS with 50,000 samples, sweeping classifier-free guidance in 0.1 increments to obtain the best FID value for each setting. CFG is applied at all scales. To account for small variations in evaluation and sampling, for each point we record values with FID within 2% (relative) of its minimum, and plot the mean, min and max IS of these with error bars.
@@
Notably, when viewed by the total number of steps (Fig. 5 top-right), scale ratios 2, 3 and 4 all achieve similar performance for each step count, even though the steps per scale differ across scale factors. The lines for these three ratios overlap tightly. However, all outperform the single-scale baseline as well as slower $\sqrt{2}$ scaling. This indicates that while multiscale conditioning is important, the exact scale factor is not very sensitive. Indeed, our checkerboard order also includes a degree of coarse-to-fine conditioning, simply due to its spatially balanced sampling, which may contribute to the insensitivity of the multiscale ratio. Thus the total number of steps in the conditional chain is the dominant performance factor.
Moreover, best performance is achieved at around 16 total steps for all of the multiscale models, while steps beyond this lead to little if any benefit. This makes sense, since partitioning our balanced checkerboard order into more steps spaces out the locations within each step, and coarse-to-fine conditioning is already included in the multiscale pyramid. Within-scale steps are needed largely to condition within the local upsampling windows. However, while between- and within-scale conditioning can model overlapping dependencies, we still find benefit to including both.
@@ 4.5 Entropy Analysis
To further illustrate the behavior of multiscale checkerboard sampling, we plot measurements of entropy for the multinomial distribution at each token during sampling.
Fig. 6 shows aggregate measurements taken over 10,000 random samples of our L model. We show mean entropy as well as shaded regions corresponding to 25th and 75th percentiles, over each sampling step. As expected, entropy decreases as steps progress and samples resolve modes and variance by conditioning. Between scales, though, there is a jump in entropy, as additional higher-resolution details are introduced, followed by continued decrease as sampling within the scale develops. Interestingly, on average, the largest entropy drop within each scale occurs half-way through the ordering, when every-other location has been filled.
Fig. 7 shows entropy for a few individual samples, both over steps, and over locations. Here, we see a trend consistent with the aggregate measures, but also substantial variation, since the token distribution depends heavily on the image content being generated. Additionally, the entropy maps clearly display checkerboard patterns, a result of our sampling order. This underscores our method’s ability to model local conditional dependencies, particularly between adjacent locations.
@@ 5 Conclusion
We have described a multiscale image generation method based on a spatially-balanced progressive checkerboard sampling order. By modeling both between-scale and within-scale conditional dependencies, our method is able to generate images with fewer sampling steps and competitive inference time compared to recent methods. Our experiments also show that in this setting, several multiscale pyramid ratios lead to similar performance for a given total number of sampling steps, enabling our method to use large scaling ratios of up to 4x. Future work may include extensions to video or text modalities, or further studying the relationship between scales and steps in the context of mutual information and dependency modeling.

### v2

@@ 4.3 Relationship Between Scale Ratio and Sampling Steps
To evaluate and compare different scale factors and sampling steps, we train our small (S) sized model using four scale-up factors: $\sqrt{2},2,3$ and $4$, as well as a single-scale baseline (exact sizes in Fig 4(b)). We then evaluate each one using different numbers of sampling steps, varying the number of steps per scale, which corresponds to partitioning using the checkerboard pattern. Note that in our implementation, each scale ratio requires retraining, but the number of sampling steps can be varied at inference.
We evaluate each condition using FID and IS with 50,000 samples, sweeping classifier-free guidance in 0.1 increments to obtain the best FID value for each setting. CFG is applied at all scales. To account for small variations in evaluation and sampling, for each point we record values with FID within 2% (relative) of its minimum, and plot the mean, min and max IS of these with error bars.
@@
However, when viewed by the total number of steps, scale ratios 2, 3 and 4 all achieve similar performance, even though each scale ratio uses a different number of steps per scale to reach the total. The lines for these three ratios all overlap tightly (Fig. 5, top-right). In addition, all outperform the single-scale baseline as well as slower $\sqrt{2}$ scaling. This indicates that while multiscale conditioning is important, the exact scale factor is not very sensitive: there are multiple ways to allocate a given number of steps among scales and obtain similar performance. Indeed, our checkerboard order also includes a degree of coarse-to-fine conditioning, simply due to its spatially balanced sampling, which may contribute to this effect. Thus the total number of steps in the conditional chain is the dominant performance factor.
Moreover, best performance is achieved at around 17 total steps for all of the multiscale models, while steps beyond this add little if any benefit. This makes sense, since partitioning our balanced checkerboard order into more steps spaces out the locations within each step, and coarse-to-fine conditioning is already included in the multiscale pyramid. Within-scale steps are needed largely to condition within the local upsampling windows. However, while between- and within-scale conditioning can model overlapping dependencies, we still find benefit to including both.
@@ 4.5 Entropy Analysis
To further illustrate the behavior of multiscale checkerboard sampling, we plot measurements of entropy for the multinomial distribution at each token during sampling. Fig. 6 shows aggregate measurements taken over 10,000 random samples of our L model. We show mean entropy as well as shaded regions corresponding to 25th and 75th percentiles, over each sampling step. As expected, entropy decreases as steps progress and samples resolve modes and variance by conditioning. Between scales, though, there is a jump in entropy, as additional higher-resolution details are introduced, followed by continued decrease as sampling within the scale develops. Interestingly, on average, the largest entropy drop within each scale occurs half-way through the ordering, when every-other location has been filled.
Fig. 7 shows entropy for individual samples, both over steps, and over locations. Here, we see a trend consistent with the aggregate measures, but also substantial variation, since the token distribution depends heavily on the image content being generated. Additionally, the entropy maps clearly display checkerboard patterns, a result of our sampling order. This underscores our method’s ability to model local conditional dependencies, particularly between adjacent locations.
@@ 5 Conclusion
We have described a multiscale image generation method based on a spatially-balanced progressive checkerboard sampling order. By modeling both between-scale and within-scale conditional dependencies, our method is able to generate images with fewer sampling steps and competitive inference time compared to recent methods. In this setting, our experiments show that several multiscale pyramid ratios lead to similar performance for a given number of total sampling steps, enabling our method to use large scaling ratios of up to 4x. Future work may include extensions to video or text modalities, or further studying the relationship between scales and steps in the context of mutual information and dependency modeling.